How long does it take to go 1km at 2.5 mph
Answer by TimothyLamb(4379)

1 km = 0.621371 mi
---
s = d / t
t = d / s
---
t = (1 km * 0.621371 mi/km ) / 2.5 mph ~= 0.2485 hours ~= 14.9 minutes
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php